a firework is luanched upward into the air. at 240ft above the ground. it burst into a shower of stars. how long does it take to reach 240ft? to find the answer, use the equation h=-16t^2+32t in which (h) is the hight above the ground at time (t) in seconds.
Answer by ankor@dixie-net.com(22740)

a firework is luanched upward into the air. at 240ft above the ground. it burst into a shower of stars. how long does it take to reach 240ft? to find the answer, use the equation h=-16t^2+32t in which (h) is the height above the ground at time (t) in seconds.
:
There is something wrong with this problem; 32t represent the velocity and that is not enough for it to ever get to 240 ft
:
Max height using the vertex (x = -b/(2a)):
x = -32/2(-16)
x = + 1 second is the vertex
:
Substitute 1 second in the original equation to find the max:
h = -16(1^1) + 32(1)
h = +16 ft is max height